OpenStudy (muscrat123):

The volume of a cone is 251.2 cubic inches. The height of the cone is 15 inches. What is the radius of the cone, rounded to the nearest inch? Use π = 3.14. 16 inches 8 inches 6 inches 4 inches

OpenStudy (muscrat123):

@dan815

OpenStudy (sepeario):

What we have to do with this question is work backwards. Do you know the formula to find the volume of a cone?

OpenStudy (muscrat123):

yes. i know that i have to work backwards, i just got an answer not an answer choice

OpenStudy (sepeario):

What did you do to get your answer?

OpenStudy (muscrat123):

idek!

OpenStudy (sepeario):

Okay let me show you how I would do it. Volume of cone = \[\frac{ 1 }{ 3 }\pi r^2 h\] (where r is radius and h is height).

OpenStudy (sepeario):

So we could just substitute the values we already have to find the values that are unknown; \[251.2=\frac{ 1 }{ 3 } 3.14 * r^2*15\]

OpenStudy (sepeario):

I'm sorry if that's a bit hard to read but do you understand so far?

OpenStudy (muscrat123):

yes

OpenStudy (muscrat123):

@Sepeario

OpenStudy (sepeario):

Then we have \[251=3.14/3 * 15 * r^2\] \[251=(3.14*15)/3 * r^2\] \[251=15.7r^2\]

OpenStudy (sepeario):

\[251/15.7=r^2\] The radius squared approximately equals 16, and the radius hence is 4.

OpenStudy (muscrat123):

so the radius is 4?

OpenStudy (sepeario):

Yes.

OpenStudy (muscrat123):

so d?